OpenStudy (anonymous):

A recipie calls for 3 eggs per serving.Eggs come in cartons of 12 .If you are going to make 14 servings how many cartons of eggs are needed?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

please explain or show work

OpenStudy (anonymous):

total eggs needed = 14*3 = 42 cartons needed = 42/12 = 3.5 so you have to buy 4 cartons